OpenStudy (anonymous):

Nitrogen has two naturally occurring isotopes, n-14 and n-15. The atomic mass of nitrogen is 14,007 amu. Which isotope is more abundant in nature? explain.

OpenStudy (australopithecus):

Create a system of equations: set x = to ratio of 14amu to 15amu abundance set y = to ratio of 15amu to 14amu abundance First equation is just what is used to calculate 14,007amu,because x and y represent their respected abundance 14amu*x + 15amu*y = 14,007 amu second equation is very simple x + y = 1 because nitrogen according to this question only has two isotopes, thus their abundance must equal 1 x + y = 1 You can think of x and y as percentages if it makes it easier, they are just divided by 100, so just say we say that 14amu had 50% prominence and 15amu had 50% prominence, 50% + 50% = 100%, if we divide these numbers by 100 and put them in ratio form (percentages just exist because people like non-decimal numbers) we get 0.5 + 0.5 = 1. Note this is just an example their prominence isn't 50/50 with these two equations, 1. solve for x in x + y = 1 2. Sub x = 1-y into 14amu*x + 15amu*y = 14,007 amu solve for y 3. sub your value of y into x + y = 1

OpenStudy (australopithecus):

4. solve for x

OpenStudy (australopithecus):

5. put both numbers into percentages by multiplying them by 100 and you have your answer